# Can Large Language Models Address Open-Target Stance Detection?

### v3

@@ Abstract
Abu Ubaida Akash  Ahmed Fahmy  Amine Trabelsi Université de Sherbrooke {akaa2803,mosa2801,amine.trabelsi}@usherbrooke.ca
@@ 4.2 Evaluation Methods
Human Evaluation: To evaluate the relevance of the generated targets and further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit) from both datasets, maintaining an equal ratio of the stance labels. Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix J. To measure inter-annotator agreement among the three raters, we compute Krippendorff’s $\alpha$ (Krippendorff, 2011), and Fleiss’s $\kappa$ (Fleiss, 1971) for each model, as shown in Appendix Table 5. We observe a solid overall agreement of $\alpha$=0.76 and $\kappa$=0.664.
@@ 5.1 Target Generation (TG)
As shown in Table 1, both human evaluation (HE) and BTSD indicate that LLMs generate higher-quality targets in explicit cases compared to TSE-BestGen, the best model for generating targets not matched to a predefined list. Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower. This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K). In contrast, LLMs mostly generate targets that are highly or partially related to golden targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models. However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 1), with the best-performing LLM getting the HE score slightly lower than $0.5$ (refers to partially related target) on average. A manual inspection of GPT-3.5’s 30 incorrect predictions suggests that the lower score in non-explicit cases is due to insufficient context, often resulting from how the TSE test set is constructed. The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix I.1).
@@ 5.2 Stance Detection (SD)
In general, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets. In some cases of TG+SD setting, though, the open LLMs achieve comparatively higher scores. Comparing approaches, LLMs tend to perform better when prompted with the holistic TG&SD, though their TG+SD results remain competitive, specifically in non-explicit cases across both datasets.
@@ Appendix E VAST Dataset Processing
Originally, the VAST dataset is a multi-target dataset, with multiple targets assigned to each sample text. However, since our work focuses on single-target analysis, we processed the VAST dataset accordingly to align it with our study (see Table 4). In VAST, there are two types of target columns: "ori_topic" which contains targets heuristically extracted, and "new_topic," which contains the final annotated targets determined by annotators. The "ori_topic" often includes duplicate targets for each text, while the "new_topic" provides unique targets with different stances.
@@ Appendix H SemSim Details
Appendix I Explanation with Samples in TG and SD
I.1 Target Generation
@@ I.2 Stance Detection
Appendix J Human Annotator Agreement
@@
Appendix K Human Annotation Score Distribution
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. The details on human annotator guideline are provided in Appendix J, and the annotator agreement is in Appendix Table 5.

### v4

@@ Abstract
Abu Ubaida Akash  Ahmed Fahmy  Amine Trabelsi Department of Computer Science, Université de Sherbrooke {akaa2803,mosa2801,amine.trabelsi}@usherbrooke.ca
@@ 4.2 Evaluation Methods
Human Evaluation: To evaluate the relevance of the generated targets and further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit) from both datasets, maintaining an equal ratio of the stance labels. Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix K. To measure inter-annotator agreement among the three raters, we compute Krippendorff’s $\alpha$ (Krippendorff, 2011), and Fleiss’s $\kappa$ (Fleiss, 1971) for each model, as shown in Appendix Table 6. We observe a solid overall agreement of $\alpha$=0.76 and $\kappa$=0.664.
@@ 5.1 Target Generation (TG)
As shown in Table 1, both human evaluation (HE) and BTSD indicate that LLMs generate higher-quality targets in explicit cases compared to TSE-BestGen, the best model for generating targets not matched to a predefined list. Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower. This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix M). In contrast, LLMs mostly generate targets that are highly or partially related to golden targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models. However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 1), with the best-performing LLM getting the HE score slightly lower than $0.5$ (refers to partially related target) on average. A manual inspection of GPT-3.5’s 30 incorrect predictions suggests that the lower score in non-explicit cases is due to insufficient context, often resulting from how the TSE test set is constructed. The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix J.1).
@@ 5.2 Stance Detection (SD)
Overall, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets. In some cases of TG+SD setting, though, the open LLMs achieve comparatively higher scores. Comparing approaches, LLMs tend to perform better when prompted with the holistic TG&SD, though their TG+SD results remain competitive, specifically in non-explicit cases across both datasets.
@@ 6 Conclusion
  • We conducted a thorough investigation of simple prompting for LLMs, inspired from the stance detection literature (§3.2), which has been shown to outperform other strategies, and provides the answers of our research questions already. However, there might be still room for improvement through a more detailed examination of different prompting techniques or task-specific pre-training. While we present some preliminary results on the chain-of-thought approach in Appendix L, further investigation is needed to fully understand its potential.
@@ Appendix E VAST Dataset Processing
Originally, the VAST dataset is a multi-target dataset, with multiple targets assigned to each sample text. However, since our work focuses on single-target analysis, we processed the VAST dataset accordingly to align it with our study (samples in Table 4). In VAST, there are two types of target columns: "ori_topic" which contains targets heuristically extracted, and "new_topic," which contains the final annotated targets determined by annotators. The "ori_topic" often includes duplicate targets for each text, while the "new_topic" provides unique targets with different stances.
@@ Appendix H SemSim Details
Appendix I OTSD with EZSTANCE
We explore the OTSD task using EZSTANCE (Zhao and Caragea, 2024), the largest zero-shot stance detection (ZSSD) dataset, to address how proprietary and open LLMs compare in both target generation (TG) and stance detection (SD) tasks when the target is explicitly or implicitly mentioned in the text. Given the open-target nature of our task, EZSTANCE was selected for its diverse range of domains and unique target set, reflecting real-world scenarios.
I.1 Dataset Processing
The dataset originally contains 47,316 Twitter samples across three stance categories (favor, against, and neutral). Since EZSTANCE features multiple targets for some texts, we adapted it into a single-target dataset (one target per text) to better align with our research by following a series of steps.
The original dataset is organized into two subtasks (target-based and domain-based), with each subtask containing two target types: claim and noun-phrase. First, we combine the samples from both target types and remove duplicates. For instances where both target types exist, we retain the noun-phrase version, as it poses a greater challenge in generating precise targets. Then, we merge the samples from both subtasks and deduplicate them based on their first occurrence. This process results in 9,462 unique samples with 6,873 distinct targets. We further divide the dataset into explicit and non-explicit cases, as outlined in Section 4.1, yielding 9,313 explicit and 149 non-explicit samples, where the target is mentioned explicitly or implicitly in the text, respectively.
I.2 Result Analysis
As shown in Table 5, LLMs face challenges in target generation for non-explicit samples compared to when the target is explicitly mentioned, based on BTSD and HE scores. When comparing proprietary and open models, consistent with our earlier findings (§5.1), no single LLM consistently outperforms the others. While GPT-4o tends to generate higher-quality targets in the TG+SD approach, Mistral surpasses other models in some TG&SD cases. Across both explicit and non-explicit samples, all models perform noticeably better using the TG&SD approach.
For stance detection, as shown in 5, similar to our findings with TSE and VAST in Section 5.2, LLMs perform better in SC when the target is explicitly mentioned. When comparing proprietary and open models, GPT-4o and Mistral both demonstrate competitive performance across approaches. Overall, LLMs consistently achieve better stance detection in the TG&SD approach.
TG+SD TG&SD
Model SS BTSD HE SC SS BTSD HE SC
Explicit
[…]
Non-explicit
GPT-3.5 0.76 38.78 0.377 36.69 0.81 41.89 0.402 38.97
[…]
Table 5: Comparison of TG and SD performance on EZSTANCE dataset (Zhao and Caragea, 2024) among the LLMs using various metrics: TG is evaluated with SS (SemSim), BTSD (%), and HE (Human Evaluation) scores, while SC (%) measures SD performance. Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
Appendix J Explanation with Samples in TG and SD
J.1 Target Generation
@@ I.2 Result Analysis
Appendix K Human Annotator Agreement
@@
Metrics TG+SD TG&SD
Ex N-Ex Ex N-Ex
BTSD-TSE 38.47 30.73 37.88 32.14
BTSD-VAST 41.10 41.54 41.34 40.38
SC-TSE 47.36 38.11 49.14 36.00
SC-VAST 48.47 46.46 49.00 45.41
Table 7: Caption goes here TG (evaluated by BTSD) and SD (evaluated by SC) performance using COT prompting on both TSE and VAST datasets in both Explicit and Non-explicit cases across the two approaches: TG+SD, TG&SD. Ex and N-Ex represent Explicit and Non-explicit cases, respectively.
Appendix L COT Initial Result
We initiate an exploration of alternative prompting techniques by employing the chain-of-thought (COT) prompting strategy for both target generation (TG) and stance detection (SD) tasks, across two approaches: TG+SD and TG&SD. As an initial phase, in Table 7, we assess target quality using the BTSD metric and evaluate stance detection performance with the SC metric, applied to both the TSE and VAST datasets for explicit and non-explicit cases.
Appendix M Human Annotation Score Distribution
After conducting the human annotation for target quality assessment, we plot the score distribution provided by the annotators, as shown in Figure 1. We observe that LLMs generally produce targets that are either partially or fully related to the golden targets, outperforming the TSE models in this regard. The targets generated by TSE-mapped are either identical to the golden targets or entirely unrelated, as TSE-mapped selects targets in its final stage (as described in Section 2) from a pre-defined list. In contrast, the targets generated by TSE-BestGen are often either partially or completely unrelated to the golden targets, indicating lower quality compared to LLMs.
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. The details on human annotator guideline are provided in Appendix K, and the annotator agreement is in Appendix Table 6.
Appendix N Vsiual Comparison of TG and SD Performance
To better illustrate the data in Table 1, we have plotted Figures 2 and 3, which represent the TG and SD performance, respectively. These figures provide a clearer visualization of the conclusions discussed in Sections 5.1 and 5.2.
Figure 2: Comparison of TG performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using various metrics: SS (Semantic Similarity), BTSD (%), and HE (Human Evaluation). The graphs are generated based on the data in Table 1.
Figure 3: Comparison of SD performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using SC (stance classification) score. The graphs are generated based on the data in Table 1.